OpenStudy (cherrilyn):

Evaluate the lim h->a of (1/h-1/a)/h-a.

OpenStudy (anonymous):

\[\lim_{h \rightarrow a}\frac{\frac{1}{h}-\frac{1}{a}}{h-a}\]=\[\lim_{h \rightarrow a}\frac{(a-h)/ah}{h-a}\]=\[\lim_{h \rightarrow a}\frac{-(h-a)/ah}{h-a}\]=\[\lim_{h \rightarrow a}-\frac{1}{ah}\]=\[-\frac{1}{a^2}\]

OpenStudy (cherrilyn):

thank you! Do you mind explaining the steps?